Fermi National Laboratory


Accelerator Update

Friday October 29

Tevatron

Bo Low Beta Separators HLS Installation - 75% complete

D0 Low Beta Separators HLS Installation - 30% complete

A0 Vacuum Upgrade - 100% complete

IPM installation - 80% complete. The Crystal Collimator has been installed

D4 Spool Replacement - 80%

D17 Separator Installation - 99% complete

D17 Separator Bake - Complete
D11 Separator Bake - Bake not needed

Shimming 412 Dipoles - 90% complete

Unrolling Magnet - 96% complete (Two magnets left)

D0 Separator unroll - 100% complete
B0 Separator unroll - 50% complete
Linac
The Linac is up and running.

Booster

Booster hit 4E12 at 70% efficiency. NTF is coming up, but no patients have been scheduled.

Antiproton Source

Debuncher Injection Area Beampipe and Quadrupole (Quad) replacement - Checkout in progress. Work on the quads is finished except for small things. Septum magnet pulse tested at full current, but need to be opened. The beam pipe installation in progress. The kicker beam pipe coating completed.
Debuncher Motorized Quads Installation - ~95% complete

AP1 Line - AP1 magnet moves complete. New VT108 installed and awaiting alignment. The vacuum window reconfiguration is complete. Vacuum pumpdown begins next week.

AP2 Line - Beam stop investigation continues. The shunt/power supply reconfiguration nearly complete. Alignment work largely complete.

SO Cord Remediation - Parts of AP50 work in progress. Completion in one to two weeks

D/A Lines - Alignment analysis in progress. Letting vacuum up for visual inspection and more pulse testing.

Beam Lines - Work in progress: AP1 magnet moves new VT108 & reconfigure vacuum windows, AP2 upstream vacuum will accessing beam stop

Main Injector/Recycler
Overall shut down progress - 85% complete

MI-30

Q307-309 Vacuum Modifications- 40% complete
Q303-305 Vacuum Modifications - 100% complete

MI-30 & 62

Correction Magnet Replacement - overall ~ 90% complete

MI-31

Electron Cooling overall - 69% complete

MI-10

Cooling Tank Repairs - ~90% complete

MiniBooNE

The horn will be tested today, October 29th.

NuMI

The horn has been successfully tested.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/1/2004   email Fermilab